A box of mass 8.0 kg slides at a speed of 10 ms−1 across a smooth level floor before it
encounters a rough patch of length 3.0 m. The frictional force on the box due to this
part of the floor is 70 N. What is the speed of the box when it leaves this rough surface?
What length of the rough surface would bring the box completely to rest?
1
Expert's answer
2018-09-06T16:46:08-0400
The deceleration of a box
a=F/m=(70 N)/(8.0 kg)=8.75 m/s^2
The distance
l=(v_f^2-v_i^2)/(2(-a))
So
v_f=√(v_i^2-2al)=√(〖10〗^2-2×8.75×3.0)=6.8 m/s
If v_f=0
l=(-v_i^2)/(2(-a))=(-〖10〗^2)/(2×(-8.75) )=5.7 m